OpenStudy (anonymous):

(5+i)^2 Complex number to standard form

OpenStudy (anonymous):

25 -1 + 10i

OpenStudy (anonymous):

24 + 10 i

OpenStudy (anonymous):

Can you solve this: (5+1)(5+i)? Use FOIL.

OpenStudy (anonymous):

24+10i

OpenStudy (anonymous):

so...it is in a + ib form

OpenStudy (anonymous):

It's not helpful to just post the answer, guys!

OpenStudy (anonymous):

which is standard form

OpenStudy (anonymous):

where a is real part and b is imaginary

OpenStudy (anonymous):

use (a=b)^2 = a^2 + 2ab + b^2

OpenStudy (anonymous):

here i^2 = -1 becoz i = sqrt -1